# DMol: A Schedule-Driven Diffusion Model for Highly Efficient and Versatile Molecule Generation

### v1

@@ 3.2 DMol Forward Noise Adding Strategy
The key idea behind our approach is as follows: at each forward diffusion step $t$, we randomly select $N(t)$ nodes and $M(t)$ edges from $\mathcal{G}^{0}$ contained in the complete subgraph induced by the selected edges (note that the absence of an edge is represented by a special label so that we are effectively dealing with a complete graph), and change their classes according to fixed state transition matrices $\textbf{Q}_{X}$ and $\textbf{Q}_{E}$, respectively, to obtain the graph $\mathcal{G}^{t}$. Here, $N(t)$ and $M(t)$ are functions of the time step index $t$. Note that we effectively run the diffusion process on randomly selected subsets of nodes and edges, which is one of the key differences when compared to DiGress.
@@
Figure 1: The forward noise addition process of DMol. At each time step $t$, we randomly selected a predetermined number of nodes and added noise only to the class labels of the nodes and a fixed number of edges in the subgraph induced by the nodes. Both selection numbers are controlled by the time index $t$. The top row depicts how the number of selected nodes $N(t)$ and edges $M(t)$ of the selected subgraph evolves with $t$ (we set $n=6,k=2,r=0.2$ and the values need to be floored to obtain integers).
@@
Figure 2: DMol diffusion relies on randomly selecting $N(t)$ nodes and $M(t)$ edges at each time step $t$ and diffusing their labels to generate a noisy graph $\mathcal{G}^{t}$. The denoising network $\phi_{\theta}$ learns to predict the denoised graph from $\mathcal{G}^{t}$. During inference, the denoised graph is combined with $q(\mathcal{\hat{G}}^{t-1}|\mathcal{\hat{G}}^{t},\mathcal{\hat{G}}^{0})$ to predict $\mathcal{\hat{G}}^{t-1}$.
@@
  Input: Graph $\mathcal{G}^{0}=(\textbf{X}^{0},\textbf{E}^{0})$, state transition matrices $\textbf{Q}_{X}\&\textbf{Q}_{E}$, node&edge selection functions $N(\cdot)\&M(\cdot)$.
@@
  Sample $\mathcal{\tilde{G}}^{t}=(\tilde{\textbf{X}}^{t},\tilde{\textbf{E}}^{t})$ from $\textbf{X}^{t}\textbf{Q}_{X}\times\textbf{E}^{t}\textbf{Q}_{E}$.
  Calculate the number of nodes and edges, $u=N(t)$ and $v=M(t)$, on which to diffuse.
@@
[…]
Additionally, our forward process constrains the class changes of edges to the subgraph induced by the nodes that underwent state change. This ensures that the sets of altered nodes and edges are codependent. Still, once the selections are made, the state changes of nodes and edges are governed by independent processes.
Next, we turn our attention to the choice of transition probability matrices. The results of DiGress (Vignac et al., 2023) have shown that making the transition probability from state $i$ to state $j$ proportional to the marginal probability of class $j$ within the training dataset results in more effective learning of the true data distribution compared to when using uniform transitions. We follow the same procedure and define the marginal probabilities of node classes in the training dataset as $\mathbf{m}_{x}\in\mathbb{R}^{1\times a}$ and those of edge classes as $\mathbf{m}_{e}\in\mathbb{R}^{1\times b}$, respectively. Hence,
$\textbf{Q}_{X}[i,j]=\begin{cases}\frac{\mathbf{m}_{x}[j]}{\sum_{\ell\neq i}\mathbf{m}_{x}[\ell]},&\text{if }i\neq j,\\ 0,&\text{if }i=j;\end{cases}$
$\textbf{Q}_{E}[i,j]=\begin{cases}\frac{\mathbf{m}_{e}[j]}{\sum_{\ell\neq i}\mathbf{m}_{e}[\ell]},&\text{if }i\neq j,\\ 0,&\text{if }i=j.\end{cases}$
During the forward process, we compute the distributions of all node and edge classes after state transitions. However, we only focus on the distributions corresponding to the selected $N(t)$ nodes and $M(t)$ edges, sampling from these distributions to update their states. We preserve the states of the remaining nodes from the previous step. Algorithm 1 and Figure 2 illustrate our forward process.
3.3 Sampling Process
During the sampling process, we first sample a random graph based on the marginal distributions of nodes and edges at $t=T$. This random graph is then used as the input to the denoising network $\phi_{\theta}$ to predict a denoised graph. Based on the prediction, we use the noise model to obtain the noisy graph at $t=T-1$. Subsequently, this noisy graph is fed into the denoising network, and the above process is repeated iteratively until a new graph is sampled.
@@ 3.3 Sampling Process
Algorithm 2 Sampling Process
  Input: Denoising Network $\phi_{\theta}$, node class distribution $\mathbf{m}_{x}^{T}=[p_{i}^{x,T}]$, edge class distribution $\mathbf{m}_{e}^{T}=[p_{k}^{e,T}]$, state transition matrices $\textbf{Q}_{X}\&\textbf{Q}_{E}$.
  Sample $\mathcal{\hat{G}}^{T}$ from $\mathbf{m}_{x}^{T}\times\mathbf{m}_{e}^{T}$.
  for $t=T$ to $1$ do
     $z\leftarrow f(\mathcal{\hat{G}}^{t},t)$DiGress structural and spectral features, as explained in (Vignac et al., 2023) Appendix B.2}.
     Calculate $u=N(t-1)$, $v=M(t-1)$.
     Randomly select $u$ nodes and $v$ edges and generate corresponding “filter” matrices $\textbf{K}_{x}$ and $\textbf{K}_{e}$.
  end for
After sampling the random graph, we use the denoising model to predict the clean graph and continue this process iteratively. At time step $t$, the input to the denoising model includes not only the noisy graph and the time $t$ but also two additional functions: $R_{x}(t)=N(t)/n$ and $R_{e}(t)=M(t)/(0.5\,n\,(n-1))$, which describe the proportion of nodes and edges in the noisy graph at time $t$ that undergo state transitions relative to the total number of nodes and edges, respectively. They indicate to the denoising network that the input noisy graph and the predicted graph should differ by $N(t)$ nodes and $M(t)$ edges, thereby helping the network make more accurate predictions. The sampling steps are listed in Algorithm  2. Figure  8 in the Appendix is an example of the sampling process from a random graph to a denoised graph.
The loss we use to train the denoising network differs from that of DiGress in so far that it contains additional penalties,
@@
$+\lambda_{2}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{X});\text{argmax}(\textbf{X})),N(t)\big{)}$
$+\lambda_{3}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{E});\text{argmax}(\textbf{E})),M(t)\big{)}.$
[…]
We also briefly comment on the equivariance property of the diffusion model and the invariance of the loss function used by DMol.
Proposition 3.1 (Equivariance).
The DMol diffusion model is permutation equivariant.
The result easily follows since our denoising model used the same graph transformer as DiGress, which is permutation equivariant.
Proposition 3.2 (Invariance of the Loss Function).
The DMol loss function is permutation invariant.
Our loss function comprises two components, a cross-entropy loss and an MSE loss.
For the cross-entropy loss terms,
@@
Since the overall cross-entropy loss is calculated by summing up the individual cross-entropy values for each node and edge, and the sum operation is permutation invariant, then the cross-entropy loss is also invariant.
We arrive at a similar conclusion for the MSE terms in $l_{MSE}(\hat{\textbf{p}}^{\mathcal{G}},\mathcal{G})$,
$\lambda_{2}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{X});\text{argmax}(\textbf{X})),N(t)\big{)}$
$+$ $\lambda_{3}\,\text{MSE}\big{(}D(\text{argmax}(\hat{\textbf{p}}^{E});\text{argmax}(\textbf{E})),M(t)\big{)}$
@@
where invariance follows based on the definition of $D(\cdot;\cdot)$ and the fact that the loss is computed separately for nodes/edges with different indices and then summed.
4 Ring Compression and Compressed DMol
@@
Inspired by JT-VAE (Jin et al., 2019), we incorporate into DMol a ring compression process, depicted in Figure 3. The process is designed to better preserve motifs in molecular graphs, which are predominantly carbon rings. Specifically, we convert a small (roughly $10$) predefined number of most frequently occurring carbon rings that can only form single bonds through their available carbons into supernodes, introducing in the process a new node class to represent the compressed rings. The compressed graph representations are directly fed into our DMol diffusion model. During the sampling process, the supernodes are converted back to their respective carbon rings.
Figure 3: In molecule graphs, frequently occurring ring structures that also allow for only single bonds with carbon atoms are compressed into supernodes during preprocessing. Each supernode comes with its own label. During DMol diffusion, supernodes are either converted into other classes of supernodes (e.g., rings can be converted into other classes of rings) or into “simple” nodes, (e.g., individual atoms), and vice versa. During the sampling process, the supernode is decoded back into its corresponding ring, and integrated into the graph by using the only possible bond structure. Since molecular graphs are sparse, it is highly unlikely to exceed two bonds per supernode.
@@
There are significant differences between our node compression approach and that of junction tree encoding in JT-VAE (Jin et al., 2019). First, the JT approach requires finding a minimum spanning tree during the encoding process. Since VAEs do not directly generate graphs, one has to perform complicated decoding. For example, each compressed molecular substructure can form many different types of bonds with different atoms in the reconstruction (i.e., reconstruction is not unique). The options are ranked according to a special scoring function and the overall process is computationally demanding. In contrast, our method simplifies the encoding and decoding processes by directly transforming only the most common carbon rings into supernodes which also allow only one or two bonding options. It is important to note that ring compression differs from node freezing since we allow different classes of rings to be converted into each other or simple atoms. The ring structures used in Compressed DMol for different datasets are available in the Appendix, Figures 456.
5 DiGress vs DMol

### v3

@@ 3.2 DMol Forward Noise Adding Strategy
The key idea behind our approach is as follows: at each forward diffusion step $t$, we randomly select $N(t)$ nodes and $M(t)$ edges from $\mathcal{G}^{0}$ contained in the complete subgraph induced by the selected nodes (note that the absence of an edge is represented by a special label so that we are effectively dealing with a complete graph), and change their classes according to fixed state transition matrices $\textbf{Q}_{X}$ and $\textbf{Q}_{E}$, respectively, to obtain the graph $\mathcal{G}^{t}$. Here, $N(t)$ and $M(t)$ are deterministic functions of the time step index $t$. Note that we effectively run the diffusion process on a determinist number of randomly selected subsets of nodes and edges, with the perturbed edges confined to the subgraph induced by the selected nodes. This effectively couples the node and edge diffusion processes.
@@
(a) At each time step $t$, we randomly select a predetermined deterministic number of nodes and add noise only to the class labels of these nodes and a fixed number of edges in the subgraph induced by the nodes. Both selection numbers are controlled by the time index $t$. The top row depicts how the number of selected nodes $N(t)$ and edges $M(t)$ of the selected subgraph evolves with $t$ (we set $n=6,k=2,r=0.2$).
@@
(b) DMol diffusion relies on randomly selecting $N(t)$ nodes and $M(t)$ edges from the subgraph induced by the $N(t)$ nodes at each time step $t$, and diffusing their labels to generate a noisy graph $\mathcal{G}^{t}$. The denoising network $\phi_{\theta}$ learns to predict the denoised graph from $\mathcal{G}^{t}$. During inference, the denoised graph is combined with $q(\mathcal{\hat{G}}^{t-1}|\mathcal{\hat{G}}^{t},\mathcal{\hat{G}}^{0})$ to predict $\mathcal{\hat{G}}^{t-1}$.
Figure 1: (a) The forward process; (b) DMol illustration.
For $N(t)$ and $M(t)$, we adopt cosine schedules, $N(t)=\lfloor(1-\alpha)\,n\rfloor$ and $M(t)=\lfloor(1-\alpha)\,N(t)\,(N(t)-1)\,0.5\,r\rfloor,$ where $\alpha=\cos^{2}(0.5\pi(t/T+c)/(1+c))$, and $n$ stands for the the number of nodes in the graph, $r$ is a hyperparameter and $c$ is a small positive constant. The total number of diffusion steps is fixed at T=knT=k\,n, where $k$ is a small hyperparameter (typically set to $1$ or $2$ so as to keep the number of diffusion steps as small as possible; details behind the selection of appropriate values of $k$ can be found in Appendix G.3). The maximum number of diffusion steps is proportional to the number of nodes $n$ (typically around $40$ for small molecule drugs). Hence, DMol requires significantly fewer diffusion steps than other methods, such as Digress (which requires $500$ or $1000$ steps), and adapts itself to the size of the graph.
Additionally, our forward process constrains the class changes of edges to the subgraph induced by the nodes that underwent a state change. This ensures that the sets of altered nodes and edges are co-dependent. Still, once the selections are made, the actual state changes of nodes and edges are governed by independent processes. We next turn our attention to the choice of the transition probability matrices. The results of DiGress Vignac et al. (2023) have shown that making the transition probability from class $i$ to class $j$ proportional to the marginal probability of class $j$ within the training dataset results in more effective learning of the true data distribution compared to when using uniform transitions. We follow the same procedure and define the marginal probabilities of node classes in the training dataset as $\mathbf{m}_{x}\in\mathbb{R}^{1\times a}$ and those of edge classes as $\mathbf{m}_{e}\in\mathbb{R}^{1\times b}$, respectively. Hence,
$\textbf{Q}_{X}[i,j]$ $=\begin{cases}\frac{\mathbf{m}_{x}[j]}{\sum_{\ell\neq i}\mathbf{m}_{x}[\ell]},&\text{if }i\neq j,\\ 0,&\text{if }i=j.\end{cases}\quad\textbf{Q}_{E}[i,j]=\begin{cases}\frac{\mathbf{m}_{e}[j]}{\sum_{\ell\neq i}\mathbf{m}_{e}[\ell]},&\text{if }i\neq j,\\ 0,&\text{if }i=j.\end{cases}$
During the forward process, we compute the distributions of all node and edge classes after state transitions. However, we only focus on the distributions corresponding to the selected $N(t)$ nodes and $M(t)$ edges, sampling from these distributions to update their states. We preserve the states of the remaining nodes from the previous step. Algorithm 1 and Figure 1(b) illustrate our forward process.
@@
Input: $\mathcal{G}^{0}=(\textbf{X}^{0},\textbf{E}^{0})$, state transition matrices $\textbf{Q}_{X}\&\textbf{Q}_{E}$, node&edge selection functions $N(\cdot)\&M(\cdot)$.
Process:
@@
 Sample $\mathcal{\tilde{G}}^{t}=(\tilde{\textbf{X}}^{t},\tilde{\textbf{E}}^{t})$ from $\textbf{X}^{0}\textbf{Q}_{X}\times\textbf{E}^{0}\textbf{Q}_{E}$.
 Calculate $u=N(t)$ and $v=M(t)$.
@@
Algorithm 2 Sampling Process
Input: Denoising Network $\phi_{\theta}$, node&edge class distribution $\mathbf{m}_{x}^{T}=[p_{i}^{x,T}]$$\mathbf{m}_{e}^{T}=[p_{k}^{e,T}]$, state transition matrices $\textbf{Q}_{X}\&\textbf{Q}_{E}$.
Process:
 Sample $\mathcal{\hat{G}}^{T}$ from $\mathbf{m}_{x}^{T}\times\mathbf{m}_{e}^{T}$.
for $t=T$ to 11 do
  $z\leftarrow f(\mathcal{\hat{G}}^{t},t)$ {Structural and spectral features}.
  Calculate u=N(t1)u=N(t-1), v=M(t1)v=M(t-1).
  Randomly select $u$ nodes and $v$ edges from the subgraph induced by the $u$ nodes and generate corresponding “filter” matrices $\textbf{K}_{x}$ and $\textbf{K}_{e}$.
end for
3.3 Denoising/Sampling Process
During the denoising process, we first sample a random graph based on the marginal distributions of nodes and edges at $t=T$. This random graph is then used as the input to the denoising network $\phi_{\theta}$ to predict a denoised graph. Based on the prediction, we use the noise model to obtain the noisy graph at t=T1t=T-1. Subsequently, this noisy graph is fed into the denoising network, and the above process is repeated iteratively until a new graph is sampled.
@@ 3.3 Denoising/Sampling Process
After sampling the random graph, we use the denoising model to predict the clean graph and continue this process iteratively. At time step $t$, the input to the denoising model includes not only the noisy graph and the time $t$ but also two additional functions: $R_{x}(t)=N(t)/n$ and $R_{e}(t)=M(t)/(0.5\,n\,(n-1))$, which describe the proportion of nodes and edges in the noisy graph at time $t$ that undergo state transitions relative to the total number of nodes and edges, respectively. They indicate to the denoising network that the input noisy graph and the predicted graph should differ by $N(t)$ nodes and $M(t)$ edges, thereby helping the network make more accurate predictions. The sampling steps are listed in Algorithm  2. Figure  8 in the Appendix is an example of the sampling process from a random graph to a denoised graph. Note that the posterior distribution used by the DMol sampling process, $q(\mathcal{\hat{G}}^{t-1}|\mathcal{\hat{G}}^{t},\phi_{\theta})$, factorizes as $q(\mathcal{\hat{G}}^{t-1}|\mathcal{\hat{G}}^{0})\,q_{\theta}(\mathcal{\hat{G}}^{0}|\mathcal{\hat{G}}^{t})$. This decomposition improves both computational efficiency and inference accuracy.
The loss we use to train the denoising network differs from that of DiGress in so far that it contains two additional penalties,
@@
$+\lambda_{2}\,\text{MSE}\big(D(\text{argmax}(\hat{\textbf{p}}^{X});\text{argmax}(\textbf{X})),N(t)\big)+\lambda_{3}\,\text{MSE}\big(D(\text{argmax}(\hat{\textbf{p}}^{E});\text{argmax}(\textbf{E})),M(t)\big).$
where $\hat{\textbf{p}}^{\mathcal{G}}=(\hat{\textbf{p}}^{X},\hat{\textbf{p}}^{E})$ denotes the predicted probability distributions, while $\textbf{x}_{i}$ and $\textbf{e}_{ij}$ represent the one-hot encodings of nodes and edges, respectively. CE stands for the cross-entropy loss, while MSE refers to the mean squared error loss. The term $\text{argmax}(\hat{\textbf{p}}^{X})$ equals the class index of highest probability for each node, while $\text{argmax}(\textbf{X})$ effectively converts the one-hot encoding of each node into its corresponding class index. $D(\cdot;\cdot)$ is equal to the number of nodes (edges) in the two arguments that have different classes. The first two terms are the standard node classification and edge classification losses. The third and fourth terms ensure that the differences in the number of nodes and edges of the predicted and the ground truth graphs, respectively, are as close as possible to the number of nodes and edges modified during the noise addition process. This is possible since $N(t),M(t)$ are deterministic - the added losses degrade performance otherwise.
We also briefly remark that it is easy to see that the DMol diffusion model and the modified loss function are permutation invariant. To see why the second claim is true, observe that our loss function comprises two components, a cross-entropy loss and an MSE loss. For the cross-entropy loss terms,
@@
Since the overall cross-entropy loss is obtained by summing up the individual cross-entropy values for each node and edge, and the sum operation is permutation invariant, then the cross-entropy loss is also invariant. We arrive at a similar conclusion for the MSE terms in $l_{MSE}(\hat{\textbf{p}}^{\mathcal{G}},\mathcal{G})$,
$\lambda_{2}\,\text{MSE}\big(D(\text{argmax}(\hat{\textbf{p}}^{X});\text{argmax}(\textbf{X})),N(t)\big)+\lambda_{3}\,\text{MSE}\big(D(\text{argmax}(\hat{\textbf{p}}^{E});\text{argmax}(\textbf{E})),M(t)\big),$
@@
where invariance follows based on the definition of $D(\cdot;\cdot)$ and the fact that the loss is computed separately for nodes/edges with different indices and then summed up.
3.4 Sampling Efficiency of DMol
To explain the sampling efficiency of DMol, we analyze the minimum time step increment $\delta t$ required to perturb exactly one node (i.e., the smallest discrete noise unit) during the diffusion process. We compare DMol with DiGress; however, it is important to note that the conclusion holds for all discrete diffusion models, including DiGress, DisCo, etc. The ratio of the required time step increments for DiGress and DMol satisfies
$\frac{\delta t_{\text{DiGress}}}{\delta t_{\text{DMol}}}=\frac{1}{\sum_{i}p^{x}_{i}(1-p^{x}_{i})}>1,$ (1)
where $p^{x}_{i}$ denotes the probability of node type $i$ (for derivations, see Appendix D.2). The inequality demonstrates that DMol requires fewer steps to perturb a single node. Since the total number of diffusion steps in DMol is solely determined by the number of nodes, the efficient addition of noise to nodes directly contributes to the overall efficiency of DMol. For further discussion of likelihood computations, noise distribution evolution, and expressivity refer to the Appendices C D.
4 Motif Compression
@@
To preserve motif compositions (e.g., carbon rings), we use a new motif compression feature, depicted in Figure 2. The process is designed to better preserve motifs in molecular graphs. Specifically, we convert a small predefined number (3153-15) of most frequently occurring subgraphs/motifs (which may differ for different training sets) that can only form single bonds with the scaffolds through their available carbons into supernodes, introducing in the process new node classes to represent the compressed motifs. The compressed graph representations are directly fed into our DMol diffusion model. During the sampling process, the supernodes are converted back to their respective motifs and integrated into the graph scaffold in a chemically valid manner.
Figure 2: Motifs (i.e., substructures occurring with high frequencies or frequencies higher than predicted by random models) that also allow for unique scaffold integration are compressed into supernodes with their own labels. During diffusion, supernodes are either converted into other classes of supernodes or into atomic nodes, and vice versa. During sampling, the supernode is decoded back into its corresponding motif.
@@
Although superficially similar to part of the JT-VAE Jin et al. (2019) approach, motif compression is significantly different. The JT approach uses not just motifs but a large number (roughly several hundreds) of submolecular structures, and requires finding a minimum spanning tree during the encoding process; furthermore, VAEs do not directly generate graphs, and one has to perform complicated decoding. The main issue is that each compressed molecular substructure can form many different types of bonds with different atoms during the reconstruction process. The options are ranked according to a special scoring function and the overall process is computationally demanding. The motifs used in DMol for different datasets are available in the Appendix, Figures 345.
5 Experiments
Table 1: Performance comparison on the QM9 dataset (see Appendix G.4 for more details).
Model V↑ V.U.↑ V.U.N.↑
GraphNVP 83.5 18.4 -
GDSS 96.0 94.6 -
GruM 99.4 85.1 22.2
DiGress 97.8 95.2 31.8
DisCo 98.2 95.6 57.5
DeFoG 97.9 95.9 62.8
DMol(Ours) 98.3 96.0 75.1
Table 2: Performance comparison on the same version of MOSES. For fair comparisons, we ran the DisCo code without any modifications, and DeFoG with $50$ diffusion steps to ensure similar computational complexity to our method (DeFoG also reports their results for $50$ diffusion steps).
Model V↑ U↑ N↑ Filters↑ FCD↓ SNN↑ Scaf↑
DiGress 84.8 100 94.5 97.2 1.18 0.55 14.6
DisCo 85.7 100 97.4 95.8 1.40 0.51 14.5
DeFoG 84.2 100 97.2 96.9 1.89 0.50 14.8
DMol (Ours) 87.8 100 100 97.8 1.12 0.58 14.8
We present next the results of running DMol on several benchmarking datasets. For experiments pertaining to other graph models (e.g., SBMs and planar graphs), the reader is referred to Appendix G.1.
Benchmarks. We compared DMol to several graph generation methods. The benchmarking models used in the experiments include GraphVAE Simonovsky and Komodakis (2018), GT-VAE Mitton et al. (2021), Set2GraphVAE Vignac and Frossard (2021), SPECTRE Martinkus et al. (2022b), GraphNVP Madhawa et al. (2019), GDSS Jo et al. (2022b), GruM Jo et al. (2024), DiGress Vignac et al. (2023), DisCo Xu et al. (2024b) and DeFoG Qin et al. (2025).
Data. We tested the performance of different generative models on the QM9 Wu et al. (2018), MOSES Polykovskiy et al. (2020), and GUACAMOL Brown et al. (2019) datasets. QM9 is a small dataset: we used 100100K molecules for training, 2020K for validation, and 1313K for testing. Both MOSES and GUACAMOL contain millions of molecules, with the number of heavy atoms inside a single molecule bounded by $26$ and $88$, respectively. We used 85% of the molecules for training, 5% for validation, and 10% for testing.
Setup. For each molecule dataset, we choose a different (relatively small) number of motifs to be converted into supernodes. The selected motif structures can be found in the Appendix A. We set the hyperparameters of DMol in both setting to $k=2$, $r=0.2$, $\lambda_{1}=5$, $\lambda_{2}=1$, $\lambda_{3}=1$. The number of training epochs is set to $500$. The experiments were conducted using NVIDIA H100 GPUs. The QM9 runs were executed on a single GPU, while the MOSES and GUACAMOL experiments were processed in parallel using $4$ GPUs.
Table 3: Classical validity, uniqueness and novelty performance comparison on GUACAMOL, for comparable generation complexity.
Model V↑ U↑ N↑ KL div↑ FCD↑
DiGress 84.5 100 99.8 93.1 68.4
DisCo 86.0 100 99.8 92.8 59.7
DeFoG 86.7 100 99.5 92.5 58.1
DMol (Ours) 86.7 100 100 94.2 69.3
Evaluation Metrics. The metrics used in the experiments include validity, uniqueness, and novelty. Validity (V) measures the proportion of generated molecules that pass basic valency checks (e.g., correspond to valid SMILES files). Uniqueness (U) quantifies the proportion of generated molecules that are distinct. Novelty (N) represents the percentage of generated molecules that do not appear in the training dataset. Furthermore, V.U. stands for the product of validity and uniqueness scores, while V.U.N. stands for the product of validity, uniqueness, and novelty.
Table 4: Biochemical performance comparison on MOSES and GUACAMOL. Here, CL=ChEMBL likeness score, SD=shingle distance, ANS=average number of diffusion steps, ST=sample time, and TT=training time. Note the significantly lower ANS, ST and $T$ times of DMol compared to DeFoG.
Dataset Model CL↑ QED↑ SD↑ ANS↓ ST(s)↓ TT(h)↓
MOSES DiGress 4.4965 0.8024 0.647 500 68.24 136
DisCo 4.1373 0.7518 0.685 500 65.36 132
DeFoG 4.0502 0.7475 0.693 50 5.81 115
DMol (Ours) 4.5033 0.8055 0.683 38 1.52 60
GUACAMOL DiGress 4.0412 0.5650 0.668 500 70.82 162
DisCo 3.9093 0.5366 0.679 500 68.76 155
DeFoG 3.8075 0.4927 0.687 50 7.62 132
DMol (Ours) 4.2231 0.5786 0.678 48 1.84 88
Since MOSES and GUACAMOL are benchmarking datasets, they introduce their own set of metrics for reporting results. In addition to V, U, and N, MOSES incorporates a filter score, the Frechet ChemNet Distance (FCD) Preuer et al. (2018), SNN, and scaffold similarity scores, while GUACAMOL uses FCD and KL divergence. The definition of these metrics are in Appendix E. The metrics mentioned above still have certain limitations. Since these metrics assess the overall quality of the generation process, but do not provide insights into the performance of individual molecules. To address this limitation, we also use the ChEMBL Likeness Score (CLscore) Bühlmann and Reymond (2020), Quantitative Estimation of Drug-likeness (QED) Bickerton et al. (2012) score, and our newly introduced Shingle distance (SD) (which quantifies the divergence between two molecular motif (shingle) distributions, see the Appendix F). The CLscore is calculated by considering how many substructures in a molecule also occur in the drug-like dataset ChEMBL Gaulton et al. (2016). QED evaluates drug-likeness by considering eight widely recognized molecular properties, such as molecular mass and polar surface area. More details are available in Appendix F. Another performance metrics used is the sample time (seconds) needed to generate a batch of $64$ molecules.
Results. The experimental results are presented in Tables 12,  3 and  4. On QM9, DMol achieves the best performance compared to other baseline models. In contrast, GruM, DiGress, DisCo, DeFoG have low novelty scores, indicating that most of the generated molecules may be duplicates or small perturbation of those already present in the training set. On MOSES and GUACAMOL, DMol demonstrates superior performance compared to other methods with respect to all metrics.
Table 4 presents a comparison of diffusion models with respect to the CL/QED scores, average number of diffusion steps, sample time, shingle distance and training time. The results demonstrate that DMol consistently outperforms DiGress, DisCo, and DeFoG across multiple metrics. DMol produces samples with excellent chemical properties as evidenced by high CL scores and QED values. DMol also has larger SD than DiGress, indicating its ability to generate more novel molecules. Although DisCo and DeFoG achieve even higher SDs, as expected, due to lack of motif preservation, this comes at the cost of reduced chemical property preservation. Additionally, DMol exhibits excellent computational efficiency, reducing the average number of diffusion steps by an order of magnitude and significantly accelerating both sampling and training time. For a detailed ablation study, please see Appendix G.2.
6 Motif Distribution Analysis
To validate that motif compression does not introduce biases towards the selected compressed motifs, we conducted a comprehensive analysis comparing motif probabilities between training and generated molecular sets on the MOSES dataset. We examine the top $30$ most frequent motifs, with the top $15$ (IDs 0140-14) structures chosen for compression into supernodes, and the remaining $15$ (IDs 152915-29) serving as a control group to assess potential distribution distortions.
Individual Motif Probabilities. For each motif, we compute the probability that a molecule contains that specific motif. The results (detailed in Appendix Table 10) demonstrate a close alignment between training and generated set distributions. The average supernode motif (IDs 0140-14) probability (IDs 0140-14) equals $0.1189$ in the training set and $0.1143$ in the generated set. For nonsupernode motifs (IDs 15-29), these average probabilities equal $0.0189$ and $0.0160$, respectively. This indicates that our approach does not create systematic bias between compressed and uncompressed substructures.
Motif Co-occurrence Patterns. We examined joint probabilities of frequent motif pairs to verify that our model maintains realistic structural relationships. For example, the co-occurrence probability of benzene (c1ccccc1) and pyridine (c1ccncc1) is $0.11$ in the training set versus $0.10$ in the generated set, and similar results hold for other pairs. This demonstrates that DMol closely preserves not only individual motif frequencies but also their co-occurrence patterns.
Node and Supernode Marginal Distributions. We compared the marginal distributions of all node and supernode classes of training and generated molecules (see Appendix Table 11). The distributions show strong alignment, confirming that DMol samples from the original data distribution without introducing biases. For instance, carbon atoms comprise $0.4897$ of the training atom set, while for the generated set this value equals $0.4940$; the most frequent supernode (c1ccccc1, benzene) appears with probabilities $0.0736$ and $0.0750$, respectively.
We attribute this strong preservation of substructure statistics to two key design choices: (1) using marginal distributions from the training set as priors for both supernode and atom sampling, and (2) enforcing that $N(t)$ nodes change at each timestep, ensuring adequate creation of generating nonmotif substructures. These results demonstrate that motif compression effectively preserves molecular motif distributions without introducing systematic bias.
7 Conditional Generation Based on Scaffolds and Docking Results
To avoid issues associated with direct conditional generation, we instead proposed a scaffold-informed pipeline. The key ideas are to cluster the training set molecules according to either their RDKit features, or embeddings generated by Mol2Vec or Grover Jaeger et al. (2018); Rong et al. (2020), and then run DMol on the samples in sufficiently large cluster groups separately (see Appendix G.6. For relevant docking results, please see Appendix I.
8 Conclusion
We present DMol, a highly computationally efficient motif-preserving graph diffusion for small molecule drug generation. DMol couples node and edge noise through a combination of deterministic and stochastic mechanisms, adapts the loss function and includes motif compression. On biochemically relevant performance metrics, such as the ChEMBL likeness and QED scores, it outperforms all other reported methods. Limitations and Future Work: Please refer to Appendix K and Appendix L.
Acknowledgments
We gratefully acknowledge the financial support and computational resources provided by Abbvie and the NSF grant CCF 24-02815, which were essential for conducting this research. We also extend our sincere thanks to Matthew S. Krafczyk from the National Center for Supercomputing Applications (NCSA) at the University of Illinois Urbana-Champaign for his valuable assistance in resolving technical simulation issues encountered during our experiments. Furthermore, we thank Jiwoong Jung and Ziheng Qi for useful discussion.
Appendix A Motif Compression
Figure 3: The $3$ selected motifs for QM9.
Figure 4: The $15$ selected motifs for MOSES.
Figure 5: The $15$ selected motifs for GUACAMOL.
Since each dataset has a different size of graph and complexity, we choose a different number of motifs, $K$ to be converted to supernodes. We select 33 motifs for QM9 3, $15$ motifs for MOSES 4, and $15$ motifs for GUACAMOL 5.
Appendix B DiGress vs DMol